Article View: datac
Article #8hardware status
From: utzoohenry
Date: Thu, 04 Feb 1982 15:47
Date: Thu, 04 Feb 1982 15:47
6 lines
563 bytes
563 bytes
Wern has switched the floppies from the 11/05 for the ones he's got on the 11/34; this "fixes" the floppy drive that has been dead for the last few days. He has also found and fixed the cable problem that inactivated the Pomz and Dunham crates, and more-or-less cured the power-supply problem that had Star Ring down. Star Ring is not yet back on Unix, this will probably happen tonight.
Technical Details
Message-ID:
<anews.Autzoo.1341>Path:
arc1-ca.pugleaf.net!utzoo!henryShow Full Headers
Complete Article Headers
Path: arc1-ca.pugleaf.net!utzoo!henry From: utzoo!henry Newsgroups: datac,hacknews Subject: hardware status Date: Thu, 04 Feb 1982 15:47:58 +0000 Message-ID: <anews.Autzoo.1341>